# Harm Amplification in Text-to-Image Models

### v1

@@ 4.1 Method 1: Distribution-Based Thresholds for Discretized Text Harm Scores
$HarmAmpThresh_{j}=\hat{y}_{j}=b_{1}j+b_{0}$ (2)
where $\text{where }y_{j}=\mu_{j}+2\cdot\sigma_{j}$
@@
This method of Distribution-Based Thresholds is best suited to situations where text and image safety classifier are not well calibrated (i.e., the output harm severity for an image harm classifier does not equate to the severity of harm outputted by the text harm classifier). Rather than comparing text and image harm scores directly which requires well-aligned classifiers, we determine whether an image amplifies harm by comparing its harm score to the relevant distribution within its corresponding text harm bucket, avoiding direct text-image harm score comparisons.
4.2 Method 2: Bucket Flip for Discretized Text and Image Scores
@@
Utilizing text and image harm safety scores similar to Method 1 (Section 4.1), the Bucket Flip method provides a more direct comparison of discretized harm categories across image and text. Consider a set of text harm scores ($T$) and image harm scores ($I$), assuming text and image harm scores have a similar scale. In cases where the scales differ, techniques such as z-score transformations can be applied to align the scales. We can then discretize $T$ and $I$ into the same $n$ even buckets ($B_{0},B_{1},\text{…},B_{n}$) where $l_{j}$ is the lower bound and $l_{j+1}$ is the upper bound of the bucket (see Equation 1).
@@
For a given (prompt, image) pair $(t,i)$, we categorize the text harm score into a specific bucket $B_{t}$ and the image harm score into a specific bucket $B_{i}$. If the image harm bucket $B_{i}$ is higher than the text harm bucket $B_{t}$, then we determine that harm amplification occurred. For example, suppose we have discretized harm scores into 10 buckets of ($0-0.1,0.1-0.2,\text{…},0.9-1.0$). Given an image harm score (i.e., $i=0.45$) and text harm score (i.e., $t=0.15$), the image score would fall into 5th bucket and the text score is in the 2nd bucket. As the image score bucket is higher than text score bucket, we determine that there was a bucket flip and harm amplification has occurred.
@@
The Bucket Flip method is most effective when both text and image harm classifiers scores are well aligned or when users do not have access to raw harm scores. In many real-world scenarios, safety classifiers often output discretized categories (i.e., low harm, medium harm, high harm) rather than raw harm scores, making the Bucket Flip method a practical approach for identifying harm amplification.
4.3 Method 3: Image-Text Co-embedding Based Harm Scores
@@
While training dedicated image or text harm classifiers enables the most direct harm amplification measurement methods, many people may not have the resources to label such data and thus may not be able to train the necessary safety classifiers. In this case, existing large pre-trained image-text models can serve as a substitute. Large pre-trained image-text models, such as CLIP Radford et al. (2021), embed images and texts such that the cosine distance between corresponding text and images is minimized in the embedding space. Taking advantage of the co-embedding space, harm amplification can be measured by how much closer a generated image’s embedding, $i$, is to a harmful concept in the embedding space than the prompt’s embedding, $p$, is. Although we cannot determine the precise embedding of harm concepts, we can embed words related to these harm concepts (e.g., for the harm sexually explicitly, we can use the words “sexualized", “porn", etc.), defining their text embeddings as $h_{0},h_{1},h_{2},...h_{k}$. These harm word embeddings are meant to approximate where harm concepts such as sexually explicit may lie in the CLIP embedding space. Harm amplification can then be quantified as $mean(cosine\_similarity(h_{k},i)-cosine\_similarity(h_{k},p))for\,all\,k$. This conceptually maps to the notion that harm amplification can be quantified by examining how much closer the image is to the approximated harm concepts compared to the prompt.
5 Experiments
In this section, we demonstrate how to employ the methods that we outlined in Section 4 on a large and varied T2I measurement dataset designed to simulate usage in real-world model deployment to develop a harm amplification criteria for each method. We then evaluate how well our methods perform by applying these criterion on an independent human annotated T2I evaluation dataset by calculating precision, recall, and f1-scores.
5.1 Measurement Dataset & Models
@@
Our measurement dataset consists of 497,157 prompts representing various demographics and prompt categories. These prompts were sampled from dogfood user data aimed at testing generative AI models. By employing this diverse and representative prompt dataset over curated adversarial or safety-specific prompts, we prioritize a broad spectrum of potential harmful representations that may be encountered in actual deployment. While curated adversarial prompts can be valuable for specific safety assessments, our method takes a broader perspective by aiming to understand how T2I models can generate harm amplifying images over a large range of prompts (from benign to borderline to harmful).
We used Stable Diffusion 2.1 Rombach et al. (2022) to generate images for our prompt measurement dataset. Four images were generated per prompt resulting in 1,988,628 total generated images. No NSFW or safety filters were applied during the image generation process. Thus, potentially harmful images were generated for our experiment and were not filtered or blocked.
@@
Safety machine learning classifiers were used to provide harm scores for sexually explicit content and violent content. Separate classifiers were used on the input text prompts and the output images with harm scores for each classifier ranging from 0 to 1. Fairness machine learning classifiers that identified probability estimates of gender presentation in images were also used with estimates for each gender (male, female, or undefined) ranging from 0 to 1. For the pre-trained image-text model, we use CLIP Radford et al. (2021) and generated embeddings for images and their corresponding prompts.
This large, varied measurement dataset served as the foundation for applying each of our methods to establish criteria for determining harm amplification in sexually explicit and violent content. Machine annotations for sexually explicit content, violent content, as well as CLIP embeddings served as inputs for our measurement techniques. After applying each of our measurement techniques to our measurement dataset, a criteria was obtained for each method:
  1. 1.
    Distribution-Based Threshold: Harm amplification occurs if the image harm is higher than a threshold for a given text harm bucket.
  2. 2.
    Bucket Flip: Harm amplification occurs if the image harm bucket is higher than the corresponding text harm bucket.
  3. 3.
    Image-Text Co-embedding: Harm amplification occurs if an image CLIP embedding is relatively closer to the harm concept embedding compared to the prompt embedding.
5.2 Evaluation Data & Metrics
To create the evaluation dataset, we sourced 742 prompt, image pairs from the Adversarial Nibbler data challenge Alicia Parish (2023). This public, crowd-sourced challenge aimed to identify failure-modes of DALL-E 2 AI (2023) and various Stable Diffusion versions (XL 1.0, 1.5, and 2.1) AI (2022). At least five raters annotated text and image safety independently as well as identified the type of harm (i.e., sexually explicit or violent) that was in the image if the image was deemed unsafe. Majority vote was used to determine the final label for each prompt harm and image harm label. We defined ground truth (if harm amplification occurred) as when the text prompt was considered safe by human annotators, but the image was deemed harmful. Ground truth for the absence of harm amplification was defined as when human annotators judged the text and image both to be safe. In addition to human labels, we obtained CLIP embeddings and machine annotations for sexually explicit and violent content for both prompts and images that served as inputs for each of our methods to infer whether harm amplification occurred. Subsequently, using the criteria established for each method (derived from our measurement data), we assessed the presence of harm amplification in the Nibbler evaluation for each harm, using human labels as ground truth.
We calculated precision, recall, and f1-scores for each method’s evaluation of harm amplification on the Nibbler dataset. This evaluation was conducted separately for sexually explicit and violence harm amplification. Furthermore, perceived gender was additionally machine annotated in the evaluation dataset to investigate potential disparities in rates of harm amplification across genders.
5.3 Experimental Results
5.3.1 Sexually Explicit & Violent Content Results
To quantify harm amplification using the Distribution-Based Threshold method (Section 4.1), we discretized the text harm scores for the measurement prompt data into 5 buckets for each harm type: sexually explicit and violence. Subsequently, we then obtained an image harm distribution from the measurement data for each of the 5 buckets corresponding to each harm type (see Appendix A, Figure 2A). The 95th percentile score, or 2 standard deviations above the mean, was then calculated for each distribution within its respective bucket. A first-degree polynomial was fitted across buckets on the 95th percentile score resulting in new fitted thresholds (Appendix A, Figure 2B) for each harm type. For our data, a first-degree linear function was most appropriate, but the type of function fitted on thresholds may vary depending on the statistics of the image data distribution. Having derived fitted image harm amplification thresholds for each text bucket for each harm type, we then assessed how well our measurement technique performed on the Nibbler evaluation dataset. Notably, we observed high precision (0.925) and recall (0.973) when evaluating sexually explicit content in the Nibbler dataset (see Table 1). However, we noted slightly lower precision (0.788) and recall (0.512) when evaluating this measurement technique on violent content (Table 1).
Using the Bucket Flip method (Section 4.2), we first calculated and recorded the mean and standard deviation of both the text harm scores and image harm scores for the measurement data for each harm type: sexually explicit and violence. We then did z-score transformations on both the text and image harm scores to transform them into similar scale. We then subsequently divided that range into 10 even buckets. For evaluation on the Nibbler dataset, we similarly applied z-score transformations to the machine annotations for each harm type using the corresponding mean and standard deviation on the text and image derived from the measurement data. We then bucketized the transformed scores with the bucket boundaries obtained from the measurement data and assessed whether the image buckets were higher than the text buckets. We observed high precision (0.896) and recall (0.956) when evaluating sexually explicit content and slightly lower precision (0.870) and recall (0.488) when evaluating on violent content.
While we expect the Image-Text Co-embedding method (Section 4.3) to perform worse at the task of identifying instances of harm amplification than using dedicated sexually explicit and violence text and image classifiers, this method is useful for those with access to limited resources. We used CLIP to encode images and prompts, and defined the sexually explicit harm concept using the words “porn," “sexual," and “nudity." We defined the violence harm concept using the words “violent," “weapons," and “injury". Utilizing more terms to define each harm concept did not have a significant effect on results (see Appendix B, Figure 5). Thus, we report results using the specified three terms for each harm concept that we mentioned above. Embedding and distance scores were calculated as described in Section 4.3 separately for the measurement data and evaluation data. Various harm amplification thresholds were applied (using distances calculated from the measurement data embeddings) to the evaluation dataset to obtain precision-recall curves (see Appendix B, Figure 6). Unlike the other methods, there is not an obvious way to define what the best harm amplification threshold is, but rather, practitioners can decide on an appropriate threshold based on their specific use case. For these results we set an operating recall of 0.6, and compute other precision and f1-scores from this number using the PR curves. As with other methods, we saw better results for sexually explicit harm amplification (precision of 0.821, recall of 0.600) than violence harm amplification (precision 0.390, recall 0.600).
@@ 5.3.1 Sexually Explicit & Violent Content Results
Distribution-Based Thresholds 0.925 0.973 0.948 0.788 0.512 0.621
Bucket Flip 0.896 0.956 0.925 0.870 0.488 0.625
Image-Text Co-embedding 0.821 0.600 0.693 0.390 0.600 0.473
@@
5.3.2 Gender Fairness
@@
As described in Related Work (Sections 2.1.2 & 2.2.1), the amplification of social stereotypes disproportionately affects marginalized communities Bianchi et al. (2023); Naik and Nushi (2023). We present quantitative measurement of harm amplification — particularly the gendered sexual objectification stereotype Fredrickson and Roberts (1997); Szymanski et al. (2011) — and show that it disproportionately affects images depicting perceived females. We recognize the dynamic and non-binary nature of gender; however, in this preliminary study, we focus on perceived binary (male/female) gender expressions due to the ease in methodology. We also acknowledge that we that cannot effectively infer people’s gender identity, and instead rely on perceived gender expression.
To show the prevalence of disparity in harm amplification across perceived gender expression in Nibbler dataset for the harm types of sexual explicit and violent content, we evaluate the efficacy of our proposed methods across the gender demographic on the Nibbler dataset. From the Nibbler dataset, for all images, both the number of people and their perceived gender expression were determined by machine classification Baruah et al. (2022). We annotate the tuples in which the number of perceived females are in strict majority to be a (perceived) female data point, and do the same for (perceived) male gender. We do not include any tuple in which, either there are no faces or in which the number of perceived females are equal to the number of perceived males.
Gender Expression Sexually Explicit Violence
# Text Safe Examples Harm Amplification Rate # Text Safe Examples Harm Amplification Rate
(Perceived) Female 173 0.80 60 0.43
(Perceived) Male 145 0.46 120 0.36
Table 2: Harm Amplification rates across perceived gender expressions on Nibbler data. Significance testing was done using z-test. Statistically significant numbers are in bold ($p<0.01$).
We calculate the rates of harm amplification across both the perceived gender expressions using the ground truth of the Nibbler dataset for both harm categories described in Section 3. Table 2 shows the the number of examples belonging to each category across the two perceived gender expressions. We find the harm amplification rate for (perceived) females is statistically significantly higher ($p<0.01$) for the category of sexually explicit harm amplification. This important observation quantifies the increased sexually explicit harm amplification present in T2I systems for the perceived female gender. Furthermore, the increased rate of harm amplification of violent content is not statistically significant ($p=0.181$).
A good method for detecting harm amplification should run equally well across all the demographic groups. The Distribution-Based Thresholds provide comparable efficacy in detection of harm amplification instances for sexually explicit content across both the perceived gender expressions (see Appendix C, Table 3). For the remaining methods, the precision in detecting over-sexualization in perceived females is higher than the precision in detecting it in perceived males.
6 Discussion & Conclusion

### v2

@@ Introduction
$HarmAmpThresh_{j}=\hat{y}_{j}=b_{1}j+b_{0},$ (2)
$where\quad y_{j}=\mu_{j}+2\cdot\sigma_{j}$
@@
As mentioned previously, this method works best when there is access to large amounts of data to calculate statistical measures from. Additionally, this method performs well when text and image safety classifier are not well calibrated (i.e., the output harm severity for an image harm classifier does not equate to the severity of harm outputted by the text harm classifier). Rather than comparing text and image harm scores directly which requires well-aligned classifiers, we determine whether an image amplifies harm by comparing its harm score to the relevant distribution within its corresponding text harm bucket, avoiding direct text-image harm score comparisons.
Method 2: Bucket Flip for Discretized Text and Image Scores
@@
Utilizing text and image harm safety scores similar to Method 1, the Bucket Flip method provides a more direct comparison of discretized harm categories across image and text. Consider a set of text harm scores ($T$) and image harm scores ($I$), assuming text and image harm scores have a similar scale. In cases where the scales differ, techniques such as z-score transformations can be applied to align the scales. We can then discretize $T$ and $I$ into the same $n$ even buckets ($B_{0},B_{1},\dots,B_{n}$) where $l_{j}$ is the lower bound and $l_{j+1}$ is the upper bound of the bucket (see Equation 1).
@@
For a given (prompt, image) pair $(t,i)$, we categorize the text harm score into a specific bucket $B_{t}$ and the image harm score into a specific bucket $B_{i}$. If the image harm bucket is higher than the text harm bucket ($B_{i}>B_{t}$), then we determine that harm amplification occurred. For example, suppose we have discretized harm scores into 10 buckets of ($0-0.1,0.1-0.2,\text{…},0.9-1.0$). Given an image harm score (i.e., $i=0.45$) and text harm score (i.e., $t=0.15$), the image score would fall into 5th bucket and the text score is in the 2nd bucket. As the image score bucket is higher than text score bucket, we determine that there was a bucket flip and harm amplification has occurred.
@@
Unlike Method 1, this method does not require any data and is most effective when both text and image harm classifiers scores are well aligned or when users do not have access to raw harm scores. In many real-world scenarios, safety classifiers often output discretized categories (i.e., low harm, medium harm, high harm) rather than raw harm scores, making the Bucket Flip method a practical approach for identifying harm amplification.
Method 3: Image-Text Co-embedding Based Harm Scores
@@
While training dedicated image or text harm classifiers enables the most direct harm amplification measurement methods, many researchers and practitioners may not have the resources to label such data and thus may not be able to train the necessary safety classifiers. In this case, existing large pre-trained image-text models can serve as a substitute. Large pre-trained image-text models, such as CLIP (Radford et al. 2021), embed images and texts such that the cosine distance between corresponding text and images is minimized in the embedding space. Taking advantage of the co-embedding space, harm amplification can be measured by how much closer a generated image’s embedding, $i$, is to a harmful concept in the embedding space than the prompt’s embedding, $p$, is. Although we cannot determine the precise embedding of harm concepts, we can embed words related to these harm concepts (e.g., for the sexually explicit harm category, we can use the words “sexual”, “porn”, etc. - see Appendix, Table 2), defining their text embeddings as $h_{0},h_{1},h_{2},...h_{k}$. These harm word embeddings are meant to approximate where harm concepts such as sexually explicit may lie in the CLIP embedding space. We can then calculate the average distance between the proximity of image embeddings and harm concepts compared to prompt embeddings and harm concepts:
$D_{harm}=\mu(cos\_similarity(h_{k},i)\\ -cos\_similarity(h_{k},p))\mid k\in\{0,1,2,...,K\}$ (3)
Harm amplification is determined to occur if $D_{harm}$ is greater than some predefined harm amplification threshold. This conceptually maps to the notion that harm amplification can be quantified by examining how much closer the image is to the approximated harm concepts compared to the prompt.
Experiments
In this section, we demonstrate how to apply our framework empirically using a combination of safety classifiers, embedding models, and measurement dataset. These methods are then evaluated on an independent human annotated dataset where f1-scores, precision, and recall are derived.
Measurement Dataset and Models
@@
To derive statistics for Method 1, we obtained a measurement dataset consisting of 497,157 prompts representing various demographics and prompt categories. These prompts were sampled from dogfood user data aimed at testing generative AI models. By employing this diverse and representative prompt dataset over curated adversarial or safety-specific prompts, we prioritize a broad spectrum of potential harmful representations that may be encountered in actual deployment. While curated adversarial prompts can be valuable for specific safety assessments, our method takes a broader perspective by aiming to understand how T2I models can generate harm amplifying images over a large range of prompts (from benign to borderline to harmful). We used Stable Diffusion 2.1 (Rombach et al. 2022) to generate images for our prompt measurement dataset. Four images were generated per prompt resulting in 1,988,628 total generated images. No NSFW or safety filters were applied during the image generation process. Thus, potentially harmful images were generated for our experiment and were not filtered or blocked.
@@
Safety machine learning classifiers were used to provide harm scores for sexually explicit content and violent content for Methods 1 and 2. Separate classifiers were used on the input text prompts and the output images with harm scores for each classifier ranging from 0 to 1. For the pre-trained image-text model used in Method 3, we chose CLIP Resnet 101 (Radford et al. 2021). In addition to safety classfiers and image-text embedding models, fairness machine learning classifiers (Baruah et al. 2022) that identified probability estimates of gender presentation in images were also used with estimates for each gender (male, female, or undefined) ranging from 0 to 1 in the evaluation data.
@@
Distribution-Based Thresholds 0.864 0.985 0.920 0.460 0.831 0.592
Bucket Flip 0.910 0.950 0.930 0.708 0.554 0.622
Image-Text Co-embedding 0.585 0.920 0.715 0.322 0.566 0.410
@@
Evaluation Data & Metrics
To create the evaluation dataset, we sourced 1125 prompt, image pairs from the Adversarial Nibbler data challenge (Quaye et al. 2024). This public, crowd-sourced challenge aimed to identify failure-modes of DALL-E 2 (Open AI 2023) and various Stable Diffusion versions (XL 1.0, 1.5, and 2.1) (Stability AI 2022). Five to six raters trained on AI safety policies annotated prompts and images separately for the presence of sexually explicit content or violent content. To derive a confidence score for sexually explicit and violence, we converted the ratings of the annotators into a proportional measure (i.e., if 3 out of 5 raters rated an image as violent, the violence score for that image would be 0.6 or 60%). The ground truth labels for harm amplification for each harm were then defined as whether the image confidence score for a given harm was greater than the corresponding text confidence score (i.e., if 80% of raters rated the image violent and only 60% rated the corresponding prompt violent, then the image amplified harm). If the image confidence score was less than or equal to the corresponding text confidence score for a given harm, then the ground truth was labeled as harm amplification not occurring. We calculated precision, recall, and f1-scores for each method’s evaluation of harm amplification on the Nibbler dataset. This evaluation was conducted separately for sexually explicit and violence harm amplification.
Furthermore, we conducted an additional analysis on the evaluation dataset to examine gender fairness. Specifically, we investigated whether harm amplification of sexually explicit or violent content disproportionately affected one perceived gender. Perceived gender was machine annotated in the evaluation dataset with significance testing performed on rates of harm amplification across genders.
Experimental Results
Evaluation of Methods in Detecting Harm Amp
To quantify harm amplification using Method 1, Distribution-Based Threshold Method, we discretized the text harm scores for the measurement prompt data into 5 buckets for each harm type: sexually explicit and violence. Subsequently, we then obtained an image harm distribution from the measurement data for each of the 5 buckets corresponding to each harm type (see Appendix, Figure 3A). The 95th percentile score, or 2 standard deviations above the mean, was then calculated for each distribution within its respective bucket. A first-degree polynomial was fitted across buckets on the 95th percentile score resulting in new fitted thresholds (Appendix, Figure 3B) for each harm type. For our data, a first-degree linear function was most appropriate, but the type of function fitted on thresholds may vary depending on the statistics of the image data distribution. Having derived fitted image harm amplification thresholds for each text bucket for each harm type, we then assessed how well our measurement technique performed on the Nibbler evaluation dataset. A precision of 0.864 and recall of 0.985 was observed when evaluating sexually explicit content in the Nibbler dataset (see Table 1). When evaluating violent harm amplification, we noted a precision of 0.460 and recall of 0.831.
Using Method 2, the Bucket Flip Method, we divided text and image harm scores obtained by safety classifiers applied on the Nibbler data into 5 even buckets. We then assessed for each prompt, image pair whether the image harm bucket was larger than the text harm bucket. This approach yielded better results than Method 1 with a precision of 0.910 and recall of 0.950 for sexually explicit harm amplification, and a precision of 0.708 and recall of 0.554 for violence harm amplification.
While we expect Method 3, Image-Text Co-embedding Method, to perform worse at the task of identifying instances of harm amplification than using dedicated sexually explicit and violence text and image classifiers, this method is useful for those with access to limited resources. We used CLIP to encode images and prompts, and defined sexually explicit harm concept using 15 words such as “porn”, “sexual”, and “nude” (see Appendix, Table 2). These concept words were sourced from internal adversarial testing and (Rando et al. 2022a) in their attempt to reverse-engineering Stable Diffusion’s NSFW filter. Adversarial testing was used to find 15 violence concept words such as “violence”, “weapons”, and “blood” (see Appendix, Table 2). Various harm amplification thresholds were applied to the evaluation dataset to obtain precision-recall curves (see Appendix, Figure 5). Unlike the other methods, there is not an obvious way to define what the best harm amplification threshold is, but rather, practitioners can decide on an appropriate threshold based on their specific use case. For these results we obtained a threshold that maximized the f1 score, and computed precision and recall from this number using the PR curves. As with other methods, we saw better results for sexually explicit harm amplification (precision of 0.585, recall of 0.920) than violence harm amplification (precision 0.322, recall 0.566).
Gender Disparities in Harm Amplification
@@
As described in the Related Works section, the amplification of social stereotypes disproportionately affects marginalized communities (Bianchi et al. 2023; Naik and Nushi 2023). In the evaluation dataset, we analyzed ground truth harm amplification rates and their disparate impacts on perceived genders. In this study, we recognize the dynamic and non-binary nature of gender, but ultimately chose to focus on perceived binary (male/female) gender expressions due to the ease in methodology. We also acknowledge that we cannot effectively infer people’s gender identity, and instead rely on perceived gender expression.
Machine annotations of perceived gender (Baruah et al. 2022) were collected for each image in the Nibbler evaluation dataset. We restricted our analysis to images in which there was a strict majority of a perceived gender. Images in which there were no faces or where the number of perceived males equaled that of perceved females were discarded. We then calculated the rates of harm amplification using the majority perceived gender. Our analysis reveal that perceived females have significantly higher rates of oversexualization ($p<0.001$, see Figure 2), in line with previous literature positing theories of gendered sexual objectification stereotypes (Fredrickson and Roberts 1997; Szymanski, Moffitt, and Carr 2011). There was no significant difference in violence harm amplification across perceived genders ($p=0.492$). We additionally show how well each method performed in detecting harm amplification across genders in the appendix section (see Table 3).
Figure 2: Difference in harm amplification rates across perceived genders in the Nibbler evaluation dataset. Images containing females were significantly more oversexualized than males whereas there was no significant difference in harm amplification of violence. $*p<0.05,\ **p<0.01,\ ***p<0.001$
Limitations
While this framework is shown to be effective in identifying harm amplification, we acknowledge the limitations of our methodologies and identify potential areas of future research. Method 1, the Distribution-Based Thresholds Method is highly sensitive to the measurement dataset’s distribution of scores. If the distribution is skewed (i.e., the dataset is predominantly either safer or more harmful), derived harm amplification thresholds will then be biased. Thus, when using this method, it is imperative to analyze and verify that the distribution of harm scores within the measurement dataset aligns with practitioners’ assumptions. As noted previously, Method 2, the Bucket Flip Method relies on the successful calibration of text and image safety classifiers. Uncalibrated classifiers can result in inaccurate identification of harm amplification. Future research should investigate how calibration methods and various types of safety classifiers affect this method. Method 3, the Image-Text Co-embedding Method, utilized the open source model CLIP to find the difference in embeddings of images and prompts in relation to harm concept words. However, CLIP has been shown to exhibit fairness issues particularly concerning demographic attributes such as gender and race (Luo et al. 2024; Wang, Zhang, and Sang 2022). A potential future direction of work could involve using a debiased version of CLIP similar to that of FairCLIP (Wang, Zhang, and Sang 2022) to investigate its efficacy in identifying harm amplification instances.
Across all methods, we see higher performance in identifying instances of harm amplification in sexually explicit content compared to violent content. Further research should investigate the factors contributing to this disparity in performance such as differences in safety classifier effectiveness, noise in human annotations, or biases within datasets used.
Discussion